OpenStudy (anonymous):

Determine the x- and y-intercepts of the linear equation. Type a numerical answer in the space provided. Type the x-intercept first, then the y-intercept second. Separate the intercepts with a comma. For example, if the x-intercept is (-10,0) and the y-intercept is (0,5), type -10,5. x - y = 3

OpenStudy (anonymous):

@iGreen

OpenStudy (texaschic101):

to find the x intercept, sub in 0 for y and solve for x x - y = 3 x - 0 = 3 x = 3.....so the x intercept is (3,0) to find the y intercept, sub in 0 for x x - y = 3 0 - y = 3 -y = 3 y = -3.....so your y intercept is (0,-3) so I suppose you would type : (3,-3)

OpenStudy (anonymous):

thanks